The relationship described below can be modeled using an equation.
Identify the variables and write an equation to solve the questions below.
number of subscriptions sold/pay per week
5/150
10/200
15/250
20/300
25/350

Sheryl has a part-time job working for a large company selling magazine subscriptions. Above is the pay scale she was given to calculate her weekly pay checks.
(a) Write and equation to calculate the pay per week given the number of subscriptions sold.
(b) What is her weekly pay if she sells 40 subscriptions?

Answer by stanbon(75887)

The relationship described below can be modeled using an equation.
Identify the variables and write an equation to solve the questions below.
number of subscriptions sold/pay per week
5/150
10/200
15/250
20/300
25/350
Sheryl has a part-time job working for a large company selling magazine subscriptions. Above is the pay scale she was given to calculate her weekly pay checks.
(a) Write and equation to calculate the pay per week given the number of subscriptions sold.
---
Use the points (5,150) and (10,200)
slope = (200-150)/(10-5) = 50/5 = 10
------------------------
intercept: 150 = 10(5) + b
b = 100
---
Equation:
Pay = 10x + 100 where x is # of subscriptions Sheryl sells.
------------------------------
(b) What is her weekly pay if she sells 40 subscriptions?
p(40) = 10*40 + 100 = $500
==================================
Cheers,
Stan H.
Reply to stanbon@comcast.net